# VideoGuide: Improving Video Diffusion Models without Training Through a Teacher’s Guide

### v2

@@ Abstract
Text-to-image (T2I) diffusion models have revolutionized visual content creation, but extending these capabilities to text-to-video (T2V) generation remains a challenge, particularly in preserving temporal consistency. Existing methods that aim to improve consistency often cause trade-offs such as reduced imaging quality and impractical computational time. To address these issues we introduce VideoGuide, a novel framework that enhances the temporal consistency of pretrained T2V models without the need for additional training or fine-tuning. Instead, VideoGuide leverages any pretrained video diffusion model (VDM) or itself as a guide during the early stages of inference, improving temporal quality by interpolating the guiding model’s denoised samples into the sampling model’s denoising process. The proposed method brings about significant improvement in temporal consistency and image fidelity, providing a cost-effective and practical solution that synergizes the strengths of various video diffusion models. Furthermore, we demonstrate prior distillation, revealing that base models can achieve enhanced text coherence by utilizing the superior data prior of the guiding model through the proposed method. Project Page: this http URL
@@ 1 Introduction
Text-to-image (T2I) diffusion models have greatly changed the way how visual content is created and distributed, enabling users to effortlessly generate creative images from detailed text descriptions. Now the AI community is looking deeper into the potential of T2I diffusion models, exploring their application to the higher dimensional field of video generation. Text-to-video (T2V) diffusion models aim to extend the capabilities of their image-based counterparts by generating coherent video sequences from text descriptions, handling both spatial and temporal dimensions simultaneously. However T2V diffusion models still show sub-standard performance regarding temporal consistency, and can lead to the generation of degraded samples. Poor temporal consistency is also the main challenge for a variety of tasks, such as creation of personalized T2V models. Hence, recent work (Wu et al., 2023; Qiu et al., 2023; Ge et al., 2024) aims to enhance various aspects of temporal quality, but suffers from problems such as degraded quality, slow inference, etc. In this work, we attend to the clear absence of a reliable method for refining the temporal quality of pretrained text-to-video (T2V) generation models, and propose a novel framework for improved generation that does not require any training or fine-tuning.
@@
Specifically, we introduce VideoGuide, a general framework that uses any pretrained video diffusion model as a guide during early steps of reverse diffusion sampling. Choice of the pretrained VDM is flexible: it is either identical to the VDM used for inference, or it is freely selected from all existing VDMs. In any case, the VDM that acts as the guide provides a consistent video trajectory by proceeding in its own denoising for a small number of steps. The guiding model’s denoised sample is then integrated into the original denoising process to guide the sample towards a direction with better temporal quality. Through interpolation, the sampling VDM is able to follow the temporal consistency of the guiding VDM to produce samples of enhanced quality. Such interpolation only needs to be involved in the first few steps of inference, but is strong enough to guide the entire denoising process towards more desirable results. Remarkably, interpolating information of the guiding model’s denoised sample has the effects of providing the base model a better noise prior, even guiding the model to create samples that were previously unreachable. VideoGuide is a versatile framework in that any pretrained video diffusion model can be used for distillation in a plug-and-play fashion. By incorporating a superior VDM as a video guide, our framework can be used to boost underperforming VDMs into state-of-the-art quality. This is particularly useful when the relatively underperforming VDM possesses unique traits unavailable for the superior VDM.
@@
In particular, we show two representative cases of how VideoGuide can be applied, with AnimateDiff (Guo et al., 2024) and LaVie (Wang et al., 2023). In AnimateDiff, a motion module is trained that can be interleaved into any pretrained T2I model. The scheme works for any personalized image diffusion model and grants easy application of controllable and extensible modules (Zhang et al., 2023; Guo et al., 2023), but not without consequences. Specifically, fixing the T2I weights limits interaction between the temporal module and generated spatial features, hence harming temporal consistency. Applying VideoGuide with an open-source state-of-the-art model without personalization capability  (Chen et al., 2024) as the guiding model, we can greatly enhance the temporal quality of AnimateDiff. This allows us to combine the best of both worlds: personalization and controllability is provided by the base model, while temporal consistency is refined by the guiding model. Likewise, LaVie is a multifaceted T2V model that offers various functions including interpolation and super-resolution in a cascaded generation framework, but shows substandard temporal consistency. Using VideoGuide, we can upgrade its temporal consistency with an external model while maintaining its multiple functions.
@@
    We provide evidence of prior distillation, in which the informative prior of guidance models can be utilized to create samples of improved text coherency.
Figure 2: Overall Pipeline. VideoGuide is a framework for enhancing temporal quality without additional training, leveraging the capabilities of any pretrained VDM. Throughout the denoising process of the sampling VDM, the guiding VDM receives an intermediate latent ${\boldsymbol{z}}_{t}$ and provides a temporally consistent sample ${\boldsymbol{z}}_{t-\tau}$ by proceeding in its own denoising for a small number of steps $\tau$. The sample ${\boldsymbol{z}}_{t-\tau}$ is then denoised and interpolated with the denoised ${\boldsymbol{z}}_{t}$ to produce a fused latent ${\boldsymbol{z}}^{\prime}_{t}$. Such interpolation only needs to take part in the first few steps of inference, and effectively guides samples towards a direction of improved temporal consistency. To further ensure model flexibility in refining high-frequency areas for better image quality, the latent ${\boldsymbol{z}}^{\prime}_{t}$ is passed through a Low-Pass Filter (LPF). Overall, VideoGuide is a straightforward addition to the original pipeline, yet it is powerful enough to significantly enhance temporal consistency without compromising imaging quality or motion smoothness.
@@ 2 Related works
The Diffusion Model. Diffusion probabilistic models (Ho et al., 2020) have achieved great success as generative models. To address the significant computational cost that arises from operating in pixel space, Latent Diffusion Models (LDMs) (Rombach et al., 2021) learn the diffusion process in latent space. LDMs utilize an encoder-decoder framework where the encoder ${\mathcal{E}}$ and the decoder ${\mathcal{D}}$ are trained together to reconstruct the input data. This training aims to satisfy the relation ${\boldsymbol{x}}={\mathcal{D}}({\boldsymbol{z}}_{0})={\mathcal{D}}({\mathcal{E}}({\boldsymbol{x}}))$, where ${\boldsymbol{z}}_{0}$ is the latent representation of the corresponding clean pixel image ${\boldsymbol{x}}$. Thus the forward diffusion process in latent space is defined as follows:
${\boldsymbol{z}}_{t}=\sqrt{\bar{\alpha}_{t}}{\boldsymbol{z}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},$ (1)
@@
where $\bar{\alpha}_{t}$ is a pre-determined noise scheduling coefficient, and ${\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\mathbf{I}})$ represents Gaussian noise sampled from a standard normal distribution. The reverse diffusion process is directed by a score-based neural network, denoted as the diffusion model ${\boldsymbol{\epsilon}}_{\theta}$, which is trained using the denoising score matching framework (Ho et al., 2020; Song et al., 2021b). The training objective for this model is formulated as follows:
$\min_{\theta}{\mathbb{E}}_{t,{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\mathbf{I}})}||{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t)||^{2}_{2}.$ (2)
@@
Following the formulation of DDIM (Song et al., 2021a), the reverse deterministic sampling from the posterior distribution $p({\boldsymbol{z}}_{t-1}|{\boldsymbol{z}}_{t},{\boldsymbol{z}}_{0})$ is given by:
${\boldsymbol{z}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}{\boldsymbol{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t)$ (3)
${\boldsymbol{z}}_{0|t}$ $=\frac{{\boldsymbol{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}$ (4)
@@
where the denoised sample at timestep $t$, denoted as ${\boldsymbol{z}}_{0|t}$, can be obtained using Tweedie’s formula.
Classifier Free Guidance (CFG). In conditional diffusion models, classifier free guidance (Ho & Salimans, 2021) enhances quality of generated samples by increasing the conditional likelihood through a weighted adjustment of the conditional distribution. Mathematically this is expressed as:
$\hat{{\boldsymbol{\epsilon}}}_{\theta}({\boldsymbol{z}}_{t},t)={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)+w[{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,{\bm{c}})-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)]$ (5)
where ${\bm{c}}$ and $\phi$ refer to the text condition and null condition, respectively, and $w$ refers to the guidance scale used during reverse sampling. To apply classifier free guidance to Eq. (3) and Eq. (4), we substitute ${\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t)$ with $\hat{{\boldsymbol{\epsilon}}}_{\theta}({\boldsymbol{z}}_{t},t)$ in both. Recent work (Chung et al., 2024) points out that using a high guidance scale $w$ (e.g., around 7.5) often results in issues such as abrupt changes and color saturation in the denoised sample ${\boldsymbol{z}}_{0|t}$ during the early timesteps of reverse sampling. To address these issues, CFG++ (Chung et al., 2024) introduces interpolation between the conditional estimate ${\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,{\bm{c}})$ and the unconditional estimate ${\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)$ using a lower guidance scale $w\in[0,1]$. Derived from score distillation sampling (SDS) (Poole et al., 2022), CFG++ replaces the renoising term $\hat{{\boldsymbol{\epsilon}}}_{\theta}({\boldsymbol{z}}_{t},t)$ into ${\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)$. In this case, Eq. (3) can be modified as below:
${\boldsymbol{z}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\boldsymbol{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{z}}_{t},t,\phi)$ (6)
Our proposed interpolation scheme operates on denoised samples for early timesteps, during which maintaining high-quality denoised samples is essential. Thus, we utilize CFG++ throughout the early stages of denoising to achieve smooth interpolation.
@@
Video Diffusion Model (VDM) & Consistent Video Generation. The Video Diffusion Model (VDM), originally proposed in ho2022video, operates the diffusion process in the video domain. Similar to LDMs, many recent VDMs (xing2023dynamicrafter; chen2023videocrafter1; he2022lvdm) are trained in the latent space to reduce computational cost. In Latent VDMs (LVDMs), a temporal layer is incorporated to facilitate frame interaction along the temporal axis during training. By modifying ${\boldsymbol{z}}_{t}$ to ${\boldsymbol{z}}_{t}^{1:N}$ in Eqs. (1)-(6), the diffusion model can be extended to the video domain. For simplicity, we will use the notation ${\boldsymbol{z}}_{t}$ to represent the latent for video generation instead of ${\boldsymbol{z}}_{t}^{1:N}$.
@@
One of the main challenges in utilizing diffusion models for video generation lies in maintaining temporal consistency. In the video domain, PYoCo (ge2024preservecorrelationnoiseprior) introduces a carefully designed progressive video noise prior to better leverage image diffusion models for video generation. However, PYoCo primarily focuses on the noise distribution during the training stage and requires extensive fine-tuning on video datasets. Recent work (qiu2023freenoise; reuse2023) also attempts to improve temporal consistency, but focuses more on long video generation and is not applicable to the basic 16 frame scenario. FreeInit (wu2023freeinit) addresses the issue of video consistency by iterative refinement of the initial noise. This method aims to resolve the training-inference discrepancy in video diffusion models by reinitializing noise with a spatio-temporal filter, ensuring the refined noise better aligns with the training distribution. While this approach enhances frame-to-frame consistency, it has a significant drawback: repeated iteration leads to the loss of fine details and imaging quality degradation. Additionally, the iterative nature of the method induces high computational costs, prolonging the generation process.
In VideoGuide, we aim to enhance video consistency without the aforementioned drawbacks. By integrating a small number of guidance steps into the original reverse sampling process, we are able to avoid image degradation while significantly reducing inference time compared to prior work. Furthermore, our method can incorporate external diffusion models to facilitate more temporally consistent video generation. This makes our approach particularly effective for models that struggle with temporal consistency but demonstrate strong performance in other areas (e.g., customizable T2I-based video models (guo2023animatediff)).

### v3

@@ Abstract
Text-to-image (T2I) diffusion models have revolutionized visual content creation, but extending these capabilities to text-to-video (T2V) generation remains a challenge, particularly in preserving temporal consistency. Existing methods that aim to improve consistency often cause trade-offs such as reduced imaging quality and impractical computational time. To address these issues we introduce VideoGuide, a novel framework that enhances the temporal consistency of pretrained T2V models without the need for additional training or fine-tuning. Instead, VideoGuide leverages any pretrained video diffusion model (VDM) or itself as a guide during the early stages of inference, improving temporal quality by interpolating the guiding model’s denoised samples into the sampling model’s denoising process. The proposed method brings about significant improvement in temporal consistency and image fidelity, providing a cost-effective and practical solution that synergizes the strengths of various video diffusion models. Furthermore, we demonstrate prior distillation, revealing that base models can achieve enhanced text coherence by utilizing the superior data prior of the guiding model through the proposed method. Project Page: https://dohunlee1.github.io/videoguide.github.io/
@@ 1 Introduction
Text-to-image (T2I) diffusion models have made significant advances in visual generation, enabling user interactive image generation with enriched text descriptions. Now the AI community is looking deeper into the potential of T2I diffusion models, exploring their application to the higher dimensional field of video generation. Text-to-video (T2V) diffusion models aim to extend the capabilities of their image-based counterparts by generating coherent video sequences from text descriptions, handling both spatial and temporal dimensions simultaneously.
Text-to-video (T2V) diffusion models often face a challenging trade-off between temporal consistency and image quality, where improvements in one frequently degrade the other. This compromise results in diminished perceived quality and negatively impacts downstream tasks, such as T2V personalization. Although recent works [wu2023freeinit, chen2024unictrl] have attempted to address aspects of temporal quality, they often do so at the expense of either visual fidelity or inference speed. In this work, we address the clear need for a robust approach that refines the temporal stability of pretrained T2V models without sacrificing image quality. We propose a novel framework that greatly improves the quality of generated samples without requiring any training or fine-tuning.
@@
Specifically, we introduce VideoGuide, a general framework that uses any pretrained video diffusion model as a guide during early steps of reverse diffusion sampling. The choice of pretrained teacher VDM is flexible; it can be freely selected from any existing VDMs, or even be the model itself. In any case, the VDM that acts as the guide provides a consistent video trajectory by proceeding in its own denoising for a small number of steps. The teacher model’s denoised sample is then interpolated with the original denoising process to guide the sample towards a direction with better temporal quality. Through interpolation, the student VDM is able to follow the temporal consistency of the teacher VDM to produce samples of enhanced quality. Such interpolation only needs to be involved in the first few steps of inference, but is strong enough to guide the entire denoising process towards more desirable results.
VideoGuide is a versatile framework that allows any pretrained video diffusion model to be used for distillation in a plug-and-play fashion. By integrating a high-performing VDM as a video guide, our framework elevates underperforming VDMs to achieve state-of-the-art quality, which is particularly useful when the student model possesses unique traits unavailable for the teacher model. Additionally, we find that interpolating the teacher model’s denoised sample provides the student model with an enhanced noise prior, guiding it to generate samples previously unattainable.
@@
In particular, we show two representative cases of how VideoGuide can be applied to combine the best of both worlds: unique functions provided by the student model and high temporal stability provided by the teacher model. In AnimateDiff [guo2023animatediff], a motion module is trained that can be interleaved into any pretrained T2I model. The scheme works for any personalized image diffusion model and grants easy application of controllable and extensible modules [zhang2023adding, guo2023sparsectrl], but not without consequences. Specifically, fixing the T2I weights limits interaction between the temporal module and generated spatial features, hence harming temporal consistency. Applying VideoGuide with an open-source state-of-the-art model without personalization capability [chen2024videocrafter2] as the teacher model, we can greatly enhance the temporal quality of AnimateDiff. Thus, personalization and controllability is provided by the student model, while temporal consistency is refined by the teacher model. Likewise, LaVie [wang2023lavie] is a multifaceted T2V model that offers various functions including interpolation and super-resolution in a cascaded generation framework, but shows substandard temporal consistency. Using VideoGuide, we can upgrade its temporal consistency with an external model while maintaining its multiple functions.
@@
    We provide evidence of prior distillation, in which the informative prior of teacher models can be utilized to create samples of improved text coherency.
Figure 2: Overall Pipeline. VideoGuide is a framework for enhancing temporal quality without additional training, leveraging the capabilities of any pretrained VDM. Throughout the denoising process of the sampling VDM, the guiding VDM receives an intermediate latent ${\bm{z}}_{t}$ and provides a temporally consistent sample ${\bm{z}}_{t-\tau}$ by proceeding in its own denoising for a small number of steps $\tau$. The sample ${\bm{z}}_{t-\tau}$ is then denoised and interpolated with the denoised ${\bm{z}}_{t}$ to produce a fused latent ${\bm{z}}^{\prime}_{t}$. Such interpolation only needs to take part in the first few steps of inference, and effectively guides samples towards a direction of improved temporal consistency. To further ensure model flexibility in refining high-frequency areas for better image quality, the latent ${\bm{z}}^{\prime}_{t}$ is passed through a Low-Pass Filter (LPF). Overall, VideoGuide is a straightforward addition to the original pipeline, yet it is powerful enough to significantly enhance temporal consistency without compromising imaging quality or motion smoothness.
@@ 2 Related Works
The Diffusion Model. Diffusion probabilistic models [ho2020denoising] have achieved great success as generative models. To address the significant computational cost that arises from operating in pixel space, Latent Diffusion Models (LDMs) [rombach2021highresolution] learn the diffusion process in latent space. LDMs utilize an encoder-decoder framework where the encoder ${\mathcal{E}}$ and the decoder ${\mathcal{D}}$ are trained together to reconstruct the input data. This training aims to satisfy the relation ${\bm{x}}={\mathcal{D}}({\bm{z}}_{0})={\mathcal{D}}({\mathcal{E}}({\bm{x}}))$, where ${\bm{z}}_{0}$ is the latent representation of the corresponding clean pixel image ${\bm{x}}$. Thus the forward diffusion process in latent space is defined as follows:
${\bm{z}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{z}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},$ (1)
@@
where $\bar{\alpha}_{t}$ is a pre-determined noise scheduling coefficient, and ${\bm{\epsilon}}\sim{\mathcal{N}}(0,{\mathbf{I}})$ represents Gaussian noise sampled from a standard normal distribution. The reverse diffusion process is directed by a score-based neural network, denoted as the diffusion model ${\bm{\epsilon}}_{\theta}$, which is trained using the denoising score matching framework [ho2020denoising, song2020score]. The training objective for this model is formulated as follows:
$\min_{\theta}{\mathbb{E}}_{t,{\bm{\epsilon}}\sim{\mathcal{N}}(0,{\mathbf{I}})}||{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t)||^{2}_{2}.$ (2)
@@
Following the formulation of DDIM [song2020denoising], the reverse deterministic sampling from the posterior distribution $p({\bm{z}}_{t-1}|{\bm{z}}_{t},{\bm{z}}_{0})$ is given by:
${\bm{z}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}{\bm{z}}_{0|t}+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t)$ (3)
${\bm{z}}_{0|t}$ $=\frac{{\bm{z}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}$ (4)
@@
where the denoised sample at timestep $t$, denoted as ${\bm{z}}_{0|t}$, can be obtained using Tweedie’s formula [efron2011tweedie].
Guidance as Optimization Problems. Applying guidance for diffusion models during sampling time assists diffusion models in exploring the latent space with fidelity to the desired manifold, yielding samples that are tailored to specific criteria. Specifically, guidance can be viewed as addressing the optimization problem $\min_{z\in M}\ell({\bm{z}})$ for a given loss function $\ell({\bm{z}})$, where $M$ represents the clean data manifold.
A direct approach can be seen in the context of diffusion model-based inverse problem solvers [chung2022improving, chung2023diffusion, chung2024decomposed]. In particular, diffusion posterior sampling (DPS) [chung2023diffusion] defines the loss function as the manifold-constrained gradient (MCG) [chung2022improving] of a noisy sample ${\bm{z}}_{t}\in M_{t}$. The loss function $\nabla_{{\bm{z}}_{t}}^{mcg}\ell({\bm{z}}_{t})=\nabla_{{\bm{z}}_{t}}\ell({\bm{z}}_{0|t})$ is then incorporated into Eq. (3) as follows:
${\bm{z}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}({\bm{z}}_{0|t}-\gamma_{t}\nabla_{{\bm{z}}_{t}}\ell({\bm{z}}_{0|t}))$ (5)
$\quad+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t)$
where $\gamma_{t}$ denotes step size. To reduce computational overhead, Eq. (5) can be equivalently viewed as follows:
${\bm{z}}_{t-1}$ $\approx\sqrt{\bar{\alpha}_{t-1}}({\bm{z}}_{0|t}-\gamma_{t}\nabla_{{\bm{z}}_{0|t}}\ell({\bm{z}}_{0|t}))$ (6)
$\quad+\sqrt{1-\bar{\alpha}_{t-1}}{\bm{\epsilon}}_{\theta}({\bm{z}}_{t},t)$
This approach, known as decomposed diffusion sampling (DDS) [chung2024decomposed], avoids computing the score Jacobian, aligning with methods from [poole2022dreamfusion].
Depending on how the loss function $\ell({\bm{z}})$ is defined, it is possible to address various tasks, such as solving inverse problems [chung2023diffusion] and generating text-conditioned images [chung2024cfg++].
In this work, we are the first to address the video consistency problem from the perspective of an optimization problem, introducing a novel function designed to enhance sample quality in the video domain. This function is integrated into the reverse diffusion process, similar to Eq. (6), to provide a simple yet effective guidance term.
@@
Video Diffusion Model & Consistent Video Generation. The Video Diffusion Model (VDM), originally proposed in [ho2022video], operates the diffusion process in the video domain. Similar to LDMs, many recent VDMs [xing2023dynamicrafter, chen2023videocrafter1, he2022lvdm] are trained in the latent space to reduce computational cost. In Latent VDMs (LVDMs), a temporal layer is incorporated to facilitate frame interaction along the temporal axis during training. By modifying ${\bm{z}}_{t}$ to ${\bm{z}}_{t}^{1:N}$ in Eqs. (1)-(6), the diffusion model can be extended to the video domain. For simplicity, we will use the notation ${\bm{z}}_{t}$ to represent the latent for video generation instead of ${\bm{z}}_{t}^{1:N}$.
One of the main challenges in utilizing diffusion models for video generation lies in maintaining temporal consistency. In the video domain, PYoCo [ge2024preservecorrelationnoiseprior] introduces a carefully designed progressive video noise prior to better leverage image diffusion models for video generation. However, PYoCo primarily focuses on the noise distribution during the training stage and requires extensive fine-tuning on video datasets. Recent studies  [chen2024unictrl, guo2024i4vgen] aim to enhance video consistency by leveraging techniques such as attention injection and pretrained text-to-image models, but either encounter image degradation issues [chen2024unictrl] or involve a lengthy pipeline to establish initial states [guo2024i4vgen]. Other approaches introduce improvements in temporal consistency but are primarily focused on long video generation, making them less suitable for the basic 16-frame scenario [qiu2023freenoise, reuse2023].
@@
FreeInit [wu2023freeinit] addresses the issue of video consistency by iterative refinement of the initial noise. This method aims to resolve the training-inference discrepancy in video diffusion models by reinitializing noise with a spatio-temporal filter, ensuring the refined noise better aligns with the training distribution. While this approach enhances frame-to-frame consistency, it has a significant drawback: repeated iteration leads to the loss of fine details and imaging quality degradation. Additionally, the iterative nature of the method induces high computational costs, prolonging the generation process.